### v1

# Adversarial KA

Abstract.
Regarding the representation theorem of Kolmogorov and Arnold (KA) as an algorithm for representing or <<expressing>> functions, we test its robustness by analyzing its ability to withstand adversarial attacks. We find KA to be robust to countable collections of continuous adversaries, but unearth a question about the equi-continuity of the outer functions that, so far, obstructs taking limits and defeating continuous groups of adversaries. This question on the regularity of the outer functions is relevant to the debate over the applicability of KA to the general theory of NNs.
1. Introduction
[…]
Much debate ensued [6, 7] over the practical implications of KA for the central questions of trainability and generalizability. While the early consensus seemed to be that KA was not of practical value in the evolving technology, later authors [8, 9] have had a more optimistic view, focusing on the philosophy of the KA proof which hints at an optimal division of labor between the shallow and deep layers on a NN: data preparation / dynamics applied to the data. KA gives a <<universal>> presentation of the data on a single hidden layer and then builds a dynamic to converge to the final layer functions ($g$) needed to represent a desired multivariate ($f$).
KA is essentially an algorithm for building a special type (shallow, one hidden layer; but highly non-linear) of NN to represent a given continuous function. In computer science, coding theory, and the theory of error correction, the ability of an algorithm to withstand adversarial noise serves as a fundamental measure of its robustness and power. This is typically the most demanding test to which algorithms are subjected. In this spirit, we investigate KA to see how much <<untapped>> expressivity still resides within the basic proof method. In our context, the adversary acts on the hidden layer and its effects must be countered by a new choice of $g$. Unlike error correcting codes, we presume knowledge of exactly how the adversary acted. This does not make the task of countering the adversary trivial, because we are not allowed to reach back and invert its effect on the hidden layer; we are only allowed to adjust the outer functions $g$.
We note that in a different context the theory of NNs has already been much influenced by efforts to find and tame adversarial examples [10]. For example, NNs which easily distinguish real world examples of cats from dogs but can often be spoofed by carefully engineered pictures that look to humans exactly like <<cat>> or <<dog>> but that the NN will misclassify. Our adversary is somewhat different; it is not acting on inputs but rather corrupting the inner workings of the NN.
Our conclusion is a bit surprising. We find in rather general circumstances that adversaries drawn from countable collections of homeomorphisms can be defeated, showing remarkable robustness. But we unearth a novel question of equi-continuity of our final functions $g$ (equi- in the adversary parameter), which we have been unable to answer. A positive answer means KA can defeat adversaries drawn from many continuous groups. So, KA’s exact level of adversarial robustness is framed as our final Questions 3.4, 3.5.
2. Background and statements
[…]
We consider only continuous functions with sup-norms, so we will omit it from the statements and definitions. Sometimes we will use the notation $C(X,Y)$ for such spaces of functions $X\to Y$. For a function $f\colon X\to Y\subset\mathbb{R}^{d}$ we denote by $f_{i}\colon X\to\mathbb{R}$ the composition of $f$ and projection on the $i$th coordinate.
The functions $I\to I^{2n+1}$ we will call tuples, since they may be considered as the tuples of $12n+$ 1 functions. For such functions we consider the maximum norm.
For fixed $\gamma_{1},\ldots,\gamma_{n}\in\mathbb{R}$, a tuple $\phi\colon I\to I^{2n+1}$, a function $g\colon\mathbb{R}\to\mathbb{R}$, and a homeomorphism $h\colon\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ define the approximator $\operatorname{Approx}(g,h)\colon I^{n}\to\mathbb{R}$ by
$\operatorname{Approx}(g,h)(x_{1},\ldots,x_{n}):=\sum_{j=1}^{2n+1}g\Bigl{(}\gamma_{1}h_{j}\bigl{(}\phi(x_{1})\bigr{)}+\ldots+\gamma_{n}h_{j}\bigl{(}\phi(x_{n})\bigr{)}\Bigr{)}.$
[…]
Figure 1. KA-style network with a homeomorphism $h$ acting on each input neuron in parallel
[…]
Figure 2. KA-style network with a homeomorphism $h$ acting on the sum of all neuron inputs simultaneously
Theorem 2.1 (proved in §3 below).
Let $n>1$ be an integer. Let $\mathcal{H}$ be a countable set of homeomorphisms $\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$. For any rationally independent $\gamma_{1},\ldots,\gamma_{n}\in\mathbb{R}$ there exists a tuple $\phi\colon I\to I^{2n+1}$ such that for any $f\colon I^{n}\to\mathbb{R}$ and any homeomorphism $h\in\mathcal{H}$ there exists a uniformly continuous function $g\colon\mathbb{R}\to\mathbb{R}$ such that $f=\operatorname{Approx}(g,h)$.
In other words, Theorem 2.1 states that we may overcome an adversary acting as in Figure 1. Note that we cannot overcome an adversary acting as in Figure 2, since such general homeomorphism could undo all the preparations of the first, <<blocky>> layer map and render the composition $\mathbb{R}^{n}\to\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ merely an inclusion into the first $n$ coordinates of $\mathbb{R}^{2n+1}$. For a specific example, one may take $n=2$ and $f(x,y)=xy$. It is easy to check that $f$ cannot be reconstructed at the four points (±$(\pm 1,\pm 1)$ even as $f(x,y)=g_{1}(x)+g_{2}(y)$. The difference between cases is that in the first case the hidden neurons receive information, it is <<scrambled>>, and then added, while in the second case the information is added and then scrambled.
Theorem 2.1 covers several interesting countable groups of adversaries. For example, $\mathcal{H}$ may consist of affine transformations defined by matrices and vectors with rational coefficients, or the affine group over number field, $AGL_{n}(\mathbb{K})$, or the group generated by homeomorphisms whose coordinate functions are polynomials over some number field.
[…]
Figure 3. Commuting of the <<extension>> from one to two variables and the affine transform $A$
To appreciate the theorem, consider the group of invertible affine transforms $A\colon\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ defined by rational matrices $(A^{i}_{j})\in\mathbb{Q}^{(2n+1)\times(2n+1)}$ and rational vectors $a\in\mathbb{Q}^{2n+1}$ so that
$A_{j}(x)=\sum_{i=1}^{2n+1}A^{i}_{j}(x_{i}+a_{i}).$
Then
$\operatorname{Approx}(g,A)=\sum_{j=1}^{2n+1}g\left(\sum_{i=1}^{2n+1}A^{i}_{j}\bigl{(}\gamma_{1}\phi_{i}(x_{1})+\ldots+\gamma_{n}\phi_{i}(x_{n})+(\gamma_{1}+\ldots+\gamma_{n})a_{i}\bigr{)}\right),$
i.e. affine adversaries commute (up to a translation) with <<extension>> from one to more ($n$) variables (see Figure 3 where it is shown explicitly for $n=2$). Theorem 2.1 implies that we may overcome an affine adversary acting on all input neurons at once, as in Figure 2. The latter does not hold for arbitrary adversaries, as we explained right after Theorem 2.1.
Note that there is one more natural way to place an adversary: it may act before the summation, but after the multiplication, i.e. the approximation would look like
$\sum_{j=1}^{2n+1}g\Bigl{(}h_{j}\bigl{(}\gamma_{1}\phi(x_{1})\bigr{)}+\ldots+h_{j}\bigl{(}\gamma_{n}\phi(x_{n})\bigr{)}\Bigr{)}.$
Obviously, for an affine adversary this is just the same function (up to a translation). For a general case, however, we do not know how to overcome such adversaries.
If there are several adversaries, then the approximation is still possible in spite of the following result.
Theorem 2.2 (proved in §3 below).
Let $n>1$ be an integer. Let $\mathcal{H}$ be a countable set of homeomorphisms $\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$. There exist tuples $\phi^{1},\ldots,\phi^{n}\colon I\to I^{2n+1}$ such that for any $f\colon I^{n}\to\mathbb{R}$ and any homeomorphisms $h^{1},\ldots,h^{n}\in\mathcal{H}$ there exists a uniformly continuous function $g\colon\mathbb{R}\to\mathbb{R}$ such that
$f(x_{1},\ldots,x_{n})=\sum_{j=1}^{2n+1}g\Bigl{(}h^{1}_{j}\bigl{(}\phi^{1}(x_{1})\bigr{)}+\ldots+h^{n}_{j}\bigl{(}\phi^{n}(x_{n})\bigr{)}\Bigr{)}\qquad\text{for any}\quad x_{1},\ldots,x_{n}\in I.$
3. Proof of Theorems 2.1 and 2.2
The proof of Theorem 2.1 we give is based on [11, 12]; see also [13]. The idea of using a Baire category argument is due to Kahane [14].
Note that in order to obtain rational independence one may take $\gamma_{1}=1$, $\gamma_{2}=\sqrt{2}$, $\gamma_{3}=\sqrt{3}$, $\gamma_{4}=\sqrt{5}$, and so on, taking $\gamma_{i}$ to be square roots of pairwise distinct prime numbers. Below we presume that these $\gamma_{i}$ are fixed (in some way).
Let $\alpha\in(0,\frac{1}{n+1})$. For any $f\colon I^{n}\to\mathbb{R}$ with $\left\lVert f\right\rVert=1$ and any homeomorphism $h\colon\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ define $U_{f,h}\subset C(I,I^{2n+1})$ to be the set of tuples $\phi\colon I\to I^{2n+1}$ such that there exists a uniformly continuous function $g^{f,h}\colon\mathbb{R}\to\mathbb{R}$ such that $\left\lVert g^{f,h}\right\rVert\leqslant\frac{1}{n+1}$ and
(1) $\left\lVert f-\operatorname{Approx}(g^{f,h},h)\right\rVert<\frac{n}{n+1}+\alpha.$
Lemma 3.1.
Let $n>1$ be an integer and $\alpha\in(0,\frac{1}{n+1})$. For any $f\colon I^{n}\to\mathbb{R}$ with $\left\lVert f\right\rVert=1$ and any homeomorphism $h\colon\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ the set $U_{f,h}$ is open dense in $C(I,I^{2n+1})$.
Proof.
We prove the lemma for $n=2$, $\gamma_{1}=1$ and $\gamma_{2}=\sqrt{2}$. The general case differs only by the choice of constants.
Since $g^{f,h}$ are uniformly continuous and (1) is an open condition, it is sufficient to prove the denseness. That is, for any $\phi^{0}\colon I\to I^{5}$ we must construct $\phi\colon I\to I^{5}$ such that $\left\lVert\phi-\phi^{0}\right\rVert<\varepsilon$ and (1) holds for some appropriate $g^{f,h}$.
Let $N$ be a large positive integer to be specified later. The union of red intervals of rank $i$, $1\leqslant i\leqslant 5$, is
$I\setminus\bigcup\limits_{\begin{subarray}{c}0\leqslant s\leqslant N\\ s\equiv i\pmod{5}\end{subarray}}\left(\frac{s}{N},\frac{s+1}{N}\right).$
These are mostly closed intervals of length $\frac{4}{N}$ with possibly shorter intervals or points near $\partial I=\{0,1\}$.
Next, we specify $\phi$ somewhat indirectly by giving the following conditions on $\phi^{h}:=h\phi$. Since $h$ is invertible this suffices [^2] One may think about $\phi^{h}$ as about original $\phi$ but in a new coordinate system, chosen by an adversary..
  1. (a)
    $\phi^{h}$ has range in $hI^{5}$;
  2. (b)
    $\left\lVert\phi^{h}-\phi^{0,h}\right\rVert\leqslant\varepsilon^{\prime}$, where $\varepsilon^{\prime}$ is sufficiently small so that this implies $\left\lVert\phi-\phi^{0}\right\rVert<\varepsilon$;
  3. (c)
    $\phi_{j}^{h}$ is constant and a rational number on each red interval of rank $j=1,\ldots,5$;
  4. (d)
    $\phi_{j}^{h}$ and $\phi_{j^{\prime}}^{h}$ assume distinct values on all red intervals, regardless of whether $j=j^{\prime}$.
It is obvious that (a) is possible. Since the range $hI^{5}$ is compact, we have that the inverse map $h^{-1}$ is uniformly continuous on $hI^{5}$. Then (b) is possible. Now lower bound $N$ by requiring (from the uniform continuity of $\phi$) that $\phi^{0,h}$ varies by at most $\varepsilon^{\prime}$ on each of the red intervals. Then (c,d) are possible since additionally $h(0,1)^{5}$ is open.
It remains to construct $g^{f,h}$. A red rectangle of rank $i$, $i=1,\ldots,5$, in $I^{2}$, is defined to be the Cartesian product of any two red intervals of rank $i$. Denote these rectangles by $R_{i,1},\ldots,R_{i,r}$ for some $r$ (probably dependent of $i$) in any order [^3] In the case of general $n\geqslant 2$ they would be red $n$-rectangular solids..
It can be easily shown that any $x\in I$ lies in red intervals of at least four different ranks. So, any $(x,y)\in I^{2}$ lies in red rectangles of at least three different ranks [^4] In the case $n\geqslant 2$ any point from $I^{n}$ lies in red $n$-rectangular solids of at least $n+1$ different ranks.. On such rectangles $R_{j,r}$
$\Phi_{j}^{h}(x,y):=\phi_{j}^{h}(x)+\sqrt{2}\phi_{j}^{h}(y)$
will assume a constant value which we simply denote as $\Phi_{j,r}^{h}$.
We must further lower bound $N$ by requiring (from the uniform continuity of the function $f$) that $\left\lvert f(x,y)-f(x^{\prime},y^{\prime})\right\rvert<\alpha$ for any pairs $(x,y)$ and $(x^{\prime},y^{\prime})$ located in the same $R_{j,r}$.
For each red rectangle $R_{j,r}$ choose some point $\ell_{j,r}\in R_{j,r}$.
Finally, define a piecewise linear $g^{f,h}\colon\mathbb{R}\to\mathbb{R}$ by the <<forced>> values $g^{f,h}(\Phi^{h}_{j,r}):=\frac{1}{3}f(\ell_{j,r})$ for each red rectangle $R_{j,r}$. This definition makes sense since all $\Phi_{j,r}^{h}$ are distinct, because $\sqrt{2}$ is irrational.
Clearly it is possible to extend $g^{f,h}$ from its <<forced>> values to ensure that $\left\lVert g^{f,h}\right\rVert\leqslant\frac{1}{3}$.
Let us verify (1). Since (for any continuous $g\colon\mathbb{R}\to\mathbb{R}$)
$\operatorname{Approx}(g,h)=\sum_{j=1}^{5}g\circ\Phi_{j}^{h},$
we need to prove that
$\left\lvert f(x,y)-\sum_{j=1}^{5}g\bigl{(}\Phi_{j}^{h}(x,y)\bigr{)}\right\rvert<\alpha+\frac{2}{3}\quad\text{for any}\quad(x,y)\in I^{2}.$
Take any $x,y\in I$. As we already know, there are at least three red rectangles $R_{j_{s},r_{s}}$, $s\in\{1,2,3\}$, containing the point $(x,y)$. For them we have $g^{f,A}(\Phi_{j_{s},r_{s}}^{h})=\frac{1}{3}f(\ell_{j_{s},r_{s}})$. These values differ from $\frac{1}{3}f(x,y)$ by less than $\frac{\alpha}{3}$. Hence
$\left\lvert\sum_{s=1}^{3}g^{f,h}\left(\Phi^{h}_{j_{s}}(x,y)\right)-f(x,y)\right\rvert<\alpha.$
The other two values $g^{f,h}\left(\Phi^{h}_{j}(x,y)\right)$ do not exceed $\frac{1}{3}$ by the absolute value. ∎
Lemma 3.2.
Let $n>1$ be an integer and $\lambda\in(\frac{2n+1}{2n+2},1)$. Let =$\mathcal{H}=\{h_{q}\}$ be a countable set of homeomorphisms $\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$. There exists a fixed $\phi\colon I\to I^{2n+1}$ so that given any $f\colon I^{n}\to\mathbb{R}$ and any $h\in\mathcal{H}$, there is uniformly continuous $g\colon\mathbb{R}\to\mathbb{R}$ with $\left\lVert g\right\rVert\leqslant\frac{1}{n+1}\left\lVert f\right\rVert$ satisfying
$\left\lVert f-\operatorname{Approx}(g,h)\right\rVert<\lambda\left\lVert f\right\rVert.$
Proof.
W.l.o.g. assume $\left\lVert f\right\rVert=1$ and let $f_{1},f_{2},\ldots$ be an infinite sequence of functions in $C(I^{n},\mathbb{R})$ all of norm $1$ and dense in the unit sphere of $C(I^{n},\mathbb{R})$. Take $\alpha:=\lambda-\frac{2n+1}{2n+2}$. The space $C(I,I^{2n+1})$ is a separable complete metric space, so by Lemma 3.1 and the Baire category theorem,
$\bigcap_{1\leqslant p,q<\infty}U_{f_{p},h_{q}}\neq\varnothing.$
(In fact, this intersection is a dense set of <<2nd category>>.) Choose $\phi$ from this intersection. For this $\phi$ choose maps $g^{f_{p},h_{q}}$ from the definition of $U_{f_{p},h_{q}}$.
Now fix any $f\colon I^{n}\to\mathbb{R}$ and any $h=h_{q}$. We need to choose the appropriate $g$. For this, choose $f_{p}$ so that $\left\lVert f-f_{p}\right\rVert<\frac{1}{2n+2}$.
Now $g:=g^{f_{p},h_{q}}$ satisfies the requirements of the lemma, since
$\left\lVert f-\operatorname{Approx}(g,h)\right\rVert\leqslant\left\lVert f-f_{p}\right\rVert+\left\lVert f_{p}-\operatorname{Approx}(g,h)\right\rVert<\frac{1}{2n+2}+\alpha+\frac{n}{n+1}=\lambda.$
Proof of Theorem 2.1.
Take any $\lambda\in(\frac{2n+1}{2n+2},1)$. Take $\phi$ from Lemma 3.2. Fix any $f\colon I^{n}\to\mathbb{R}$ and $h\in\mathcal{H}$. Lemma 3.2 fuels a non-linear recursion (with $\phi$ fixed) in which at each step the sup-norm difference between $f$ and its $(g,h)$-approximation is decreased by a factor of $\lambda$.
Formally, let $g_{0}:=0$ and let $g_{m+1}$ be obtained by applying Lemma 3.2 for fixed $\phi$ to $f-\sum\limits_{k=0}^{m}\operatorname{Approx}(g_{k},h)$ and $h$. Hence for any integer $m>0$
$\left\lVert f-\sum\limits_{k=0}^{m}\operatorname{Approx}(g_{k},h)\right\rVert<\lambda\left\lVert f-\sum\limits_{k=0}^{m-1}\operatorname{Approx}(g_{k},h)\right\rVert<\ldots<\lambda^{m}\left\lVert f\right\rVert.$
Since $\left\lVert g_{m}\right\rVert\leqslant\frac{1}{n+1}\left\lVert f-\sum\limits_{k=0}^{m-1}\operatorname{Approx}(g_{k},h)\right\rVert\leqslant\frac{1}{n+1}\lambda^{m-1}\left\lVert f\right\rVert$, the functional series $\sum\limits_{m=0}^{\infty}g_{m}$ converges uniformly on $\mathbb{R}$ to a function $g\colon\mathbb{R}\to\mathbb{R}$. Then
$\sum_{m=0}^{\infty}\operatorname{Approx}(g_{m},h)=\operatorname{Approx}\left(\sum_{m=0}^{\infty}g_{m},h\right)=\operatorname{Approx}(g,h).$
Hence $\left\lVert f-\sum\limits_{m=0}^{\infty}\operatorname{Approx}(g_{m},h)\right\rVert=0$, and so $f=\operatorname{Approx}(g,h)$. ∎
Now we are going to give the proof of Theorem 2.2. The proof is analogous to the proof of Theorem 2.1, so we give only the sketch.
Let $\alpha\in(0,\frac{1}{n+1})$. For any $f\colon I^{n}\to\mathbb{R}$ with $\left\lVert f\right\rVert=1$ and any homeomorphisms $h^{1},\ldots,h^{n}\colon\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ define $U_{f,\{h^{i}\}}\subset C(I,I^{2n+1})^{n}$ to be the set of tuples $(\phi^{1}\colon I\to I^{2n+1},\ldots,\phi^{n}\colon I\to I^{2n+1})$ such that there exists a uniformly continuous function $g^{f,\{h^{i}\}}\colon\mathbb{R}\to\mathbb{R}$ such that $\left\lVert g^{f,\{h^{i}\}}\right\rVert\leqslant\frac{1}{n+1}$ and
(2) $\left\lvert f(x_{1},\ldots,x_{n})-\sum_{j=1}^{2n+1}g^{f,\{h^{i}\}}\Bigl{(}h^{1}_{j}\bigl{(}\phi^{1}(x_{1})\bigr{)}+\ldots+h^{n}_{j}\bigl{(}\phi^{n}(x_{n})\bigr{)}\Bigr{)}\right\rvert<\frac{n}{n+1}+\alpha\qquad\text{for any}\quad x_{1},\ldots,x_{n}\in I.$
Lemma 3.3.
Let $n>1$ be an integer and $\alpha\in(0,\frac{1}{n+1})$. For any $f\colon I^{n}\to\mathbb{R}$ with $\left\lVert f\right\rVert=1$ and any homeomorphisms $h^{1},\ldots,h^{n}\colon\mathbb{R}^{2n+1}\to\mathbb{R}^{2n+1}$ the set $U_{f,\{h^{i}\}}$ is open dense in $C(I,I^{2n+1})^{n}$.
Proof.
Since $g^{f,\{h^{i}\}}$ are uniformly continuous and (2) is an open condition, it is sufficient to prove the denseness. That is, for any $\phi^{0,1},\ldots,\phi^{0,n}\colon I\to I^{5}$ we must construct $\phi^{1},\ldots,\phi^{n}\colon I\to I^{5}$ such that $\left\lVert\phi^{i}-\phi^{0,i}\right\rVert<\varepsilon$ for each $i$ and (2) holds for some appropriate $g^{f,\{h^{i}\}}$.
The proof differs from the proof of Lemma 3.1 only in the construction of functions $\phi^{i}$. Take red intervals from the proof of Lemma 3.1. Take $\gamma_{i}$ to be square roots of distinct prime numbers. Now we specify $\phi^{i}$ somewhat indirectly by giving the following conditions on $\phi^{h^{i}}:=h^{i}\phi^{i}$. Since $h^{i}$ are invertible this suffices.
  1. (a’)
    $\phi^{h^{i}}$ has range in $h^{i}I^{5}$ for each $i=1,\ldots,n$;
  2. (b’)
    $\left\lVert\phi^{h^{i}}-\phi^{0,h^{i}}\right\rVert\leqslant\varepsilon^{\prime}$, where $\varepsilon^{\prime}$ is sufficiently small so that this implies $\left\lVert\phi^{i}-\phi^{0,i}\right\rVert<\varepsilon$;
  3. (c’)
    $\phi_{j}^{h^{i}}$ is constant and a number of the kind $q\gamma_{i}$ on each red interval of rank $j=1,\ldots,2n+1$, where $q\in\mathbb{Q}\setminus\{0\}$;
  4. (d’)
    $\phi_{j}^{h^{i}}$ and $\phi_{j^{\prime}}^{h^{i}}$ assume distinct values on all red intervals, regardless of whether $j=j^{\prime}$.
Here we need to lower bound $N$ by requiring (from the uniform continuity of $\phi^{i}$) that $\phi^{0,h^{i}}$ varies by at most $\varepsilon^{\prime}$ on each of the red intervals for any $i=1,\ldots,n$.
Then denote
$\Phi^{\{h^{i}\}}(x_{1},\ldots,x_{n}):=\phi^{h^{1}}(x_{1})+\ldots+\phi^{h^{n}}(x_{n})$
and repeat the construction of $g$ from the proof of Lemma 2.1. ∎
After this lemma, construct the universal tuples $\phi_{1},\ldots,\phi^{n}$ and approximations for each $f$ as in Lemma 3.2. Finally, apply the obtained analogue of Lemma 3.2 to fuel a non-linear recursion and thus prove Theorem 2.2.
We leave the reader with the following open questions.
Question 3.4.
Do the analogues of Theorems 2.1, 2.2 hold when the set of adversaries is the entire homeomorphism group of $\mathbb{R}^{2n+1}$?
Question 3.5.
Do the analogues of Theorems 2.1, 2.2 hold for adversaries acting after the summation, as diagramed in Figure 2, for any interesting class of adversaries beyond the cases of: 1. Countable subsets of the affine group, and 2. The continuous group of translations $\{T\}$ as discussed in the final paragraph of §4.Conclusions (below)? Both countable and continuous families of $C^{2}$-diffeomorphisms look like an interesting case to consider.
4. Conclusions
All proofs of KA using the Baire category theorem, including ours, easily pass from a countable dense collection $\{f\}$ of functions for which a <<working>> $\phi$ has been found to all functions but taking a limit. It may therefore come as a surprise that we are not able to similarly take a limit to pass from a countable collection $\mathcal{H}$ of adversaries to its closure. This would be the hoped-for conclusion. What goes wrong? The problem is that our rule for producing the outer function $g^{f,h}$ depends delicately on the location of the <<forced>> values on the real line. The function $g^{f,h}$ at these values (both in the approximation step and after convergence) bounces around between the various $\frac{1}{3}f(\ell_{j,r})$, which initially means within $[\,-\frac{1}{3},\frac{1}{3}\,]$. If such large transitions occur at near and nearer <<forced>> values the Lipshitz constant of $g^{f,h}$ will diverge. As the adversary $h$ varies (think of rotating the plane and projecting) the <<forced>> values may cross each other. This means that the functions $\{g^{f,h}\}$ are not necessarily equi-continuous w.r.t. $h$.
To give a very simple instance of this problem consider an adversary which acts on $\mathbb{R}^{5}$ merely by Euclidian translation. So, $\mathbb{R}^{5}=\{T\}$ is (also) our group of allowed adversaries. If one translates any given coordinate axis, the <<forced>> values will not cross and, trivially, the adversary may be defeated by simply pre-composing $\{g^{T}_{1},\ldots,g^{T}_{5}\}$ by the inverse translation $T^{-1}$ — if we are allowed five independent outer functions, one for each hidden layer neuron. If we continue, though, to maintain the initial rules of our game, and ask for a single outer function $g^{T}$, then the independent axial translations of the <<forced>> values, now projected onto a single Real axis, will in general cross, causing a divergence of $\mathop{\text{Lip}}(g^{T})$ and a loss of equi-continuity. This is the simplest context in which we do not know how to answer our question. As explained in our discussion of affine adversaries, translations $\{T\}$ it makes no difference whether the adversarial action occurs before or after summation is taken at the hidden layer.